A random variable x has a mean of 100 and a standard deviation of 10 Find P (85≤ x ≤ 115 )

Answer by stanbon(75887)

A random variable x has a mean of 100 and a standard deviation of 10 Find P (85≤ x ≤ 115 )
------------
z(85) = (85-100)/10 = -15/10 = -3/2
----
z(115) = (115-100)/10 = 15/10 = 3/2
----------
P(85<= x <=115) = P(-3/2 <= z <= 3/2) = normalcdf(-1.5,1.5) = 0.8664
---------------
Cheers,
Stan H.
---------------